a)     Wires manufactured for a certain computer system are specified to have a resistance of between 0.11 and 0.16 ohm. The actual measured resistances of the wires produced by Company A have a normal probability distribution, with a mean of 0.14 ohms, and a standard deviation of 0.003 ohms. What is the probability that a randomly selected wire from Company A’s production lot will meet the specifications?



1
Expert's answer
2021-07-06T15:21:10-0400

Let "X=" resistances of the wire: "X\sim N(\mu, \sigma^2)."

Given "\mu=0.14\ ohms, \sigma=0.003\ ohms."


"P(0.11<X<0.16)=P(X<0.16)-P(X\leq0.11)"

"=P(Z<\dfrac{0.16-0.14}{0.003})-P(Z\leq \dfrac{0.11-0.14}{0.003})"

"\approx P(Z<6.6667)-P(Z\leq -10)"

"\approx1"